OpenStudy (anonymous):

The principle of conservation of heat energy states that that when an object is at constant temperature or is in thermal equilibrium, it is losing and gaining heat at equal rates. the heat lost by a hot body is equal to the heat gained by the cold body in any system provided there is no heat exchange between the substances involved and their surrounding It is defined as the process in which molecules move from area of high concentration to another area of low concentration until an equilibrium concentration is established within the system under consideration (P+aV2)(V−b)=RT

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

you are referring to the chemical potential

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

so, which of the option is right?

OpenStudy (michele_laino):

using statistical physics, we can show that particles tend to move from a region of higher chemical potential, to a region of lower chemical potential, as the system approaches equilibrium

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

so at equilibrium, the chemical potential has to be equal for all subsystems of the original system Furthermore, an equilibrium state, is characterized by a definite value of Temperature, Pressure and Volume, since it has to be represented on a plane P-V for example

OpenStudy (anonymous):

so, option d

OpenStudy (michele_laino):

I think b.

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Hooke’s law states that

OpenStudy (anonymous):

the force applied on elastic material is directly proportional to the acceleration produced provided the elastic limit is not exceeded. the force applied on elastic material is directly proportional to the extension produced provided the elastic limit is exceeded. the escape thrust applied on elastic material is directly proportional to the extension produced provided the elastic limit is not exceeded. the force applied on elastic material is directly proportional to the extension produced provided the elastic limit is not exceeded

OpenStudy (michele_laino):

please wait a moment

OpenStudy (anonymous):

ok waiting

OpenStudy (michele_laino):

here the equation of a spring is: \[F = - kx\] where k is the constant spring and x is the elongation of the spring

OpenStudy (anonymous):

ok. D then

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

The latent heat of fusion of pure water is 334kJ/kg . How much energy would it take to melt 3 kg of ice at 0oC to water at 0oC 111 kJ 334 kJ 668 kJ 1 000 kJ

OpenStudy (michele_laino):

it is very simple, by definition of latent heat, the requested amount of heat is: Q=334 * 3=...?

OpenStudy (anonymous):

1002

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

so what is the right option?

OpenStudy (anonymous):

1000

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

or 668?

OpenStudy (michele_laino):

no, I don't think, since the formula, is: \[Q = \lambda \times m\] where \lambda is the latent heat = 79.7 kCal, and m is the mass of water

OpenStudy (anonymous):

ok. please continue

OpenStudy (michele_laino):

now we can write: 79.7 kCal = 79.7*4184 =333.5 kJoules so: Q= 333.5 * 3=1000.5 kJoules

OpenStudy (anonymous):

waw. u must be a genus

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

The absolute zero temperature refers to the temperature at which pure ice, water and water vapour at normal atmospheric pressure are in equilibrium theoretically all thermal motions will cease pure ice melts at normal atmospheric pressure pure ice ecomes steam at atmospheric pressure

OpenStudy (michele_laino):

sorry the unit of measure of latent heat is kJoules/Kg, or kCal/Kg

OpenStudy (michele_laino):

when T=0 K all molecules and atoms and quarks are theoretically at rest. We say theoretically, because we can not have a particle at rest since it is against the Heisenberg Principle

OpenStudy (anonymous):

so A?

OpenStudy (michele_laino):

no, A. it is not, since that is the definition of the triple point of water

OpenStudy (anonymous):

so what about B. it must be B right

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Two bodies may be said to be in thermal equilibrium if the bodies are thermally insulated from one another the bodies are not in thermal equlibrium with another body if one body loses heat to the other if there not net flow of heat between the two bodies two bodies in thermal contact

OpenStudy (anonymous):

i think D

OpenStudy (michele_laino):

yes! I think so!

OpenStudy (anonymous):

Thermal expansion of a solid material depends on the following EXCEPT the nature of the material making up the solid the range of the temperature change the initial dimensions of the solid average translational motion of constituent atoms of the material

OpenStudy (michele_laino):

for example, the thermal expantion of a rod with initial length L0, and thermal dilation coefficient \lambda, is: \[L = {L_0}(1 + \lambda \times \Delta t)\] where L is the final length, and Delta t is the change of temperature. Of course \lambda depends on the material

OpenStudy (michele_laino):

oops..expansion*

OpenStudy (anonymous):

so what is the option? D?

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

or C?

OpenStudy (michele_laino):

I think D

OpenStudy (anonymous):

Tensile strain is mathematically expressed as: Force/Area initial length/extension extension/initial lenght Stress + initial lenght

OpenStudy (michele_laino):

since L_0 is the initial length

OpenStudy (michele_laino):

by definition it is the ratio between a force and an area

OpenStudy (anonymous):

thought as much

OpenStudy (anonymous):

Which of the following does NOT determine the amount of internal energy an object has? temperature amount of material type of material shape of the object

OpenStudy (anonymous):

D?

OpenStudy (michele_laino):

yes! of course!

OpenStudy (anonymous):

A certain resistance thermometer at triple point of water has resistance of 152.0Ω . What is the temperature of the system in degrees celcius when the resistance of the thermometer is 230.51Ω

OpenStudy (anonymous):

414.2o C 141.0o C 253.2o C 80.4o C

OpenStudy (michele_laino):

here we have to apply this formula: \[\Large R\left( t \right) = {R_0}\left( {1 + \alpha t} \right)\] where \alpha is the temperatur coefficient, and R_0 is the initial resistance

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

do you know what is the value of \alpha?

OpenStudy (anonymous):

no, please what is the value?

OpenStudy (michele_laino):

wait a moment, I think that we have to apply another formula

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

please wait a moment another student asked me to help him

OpenStudy (anonymous):

ok. am waiting sir

OpenStudy (anonymous):

hello sir, still waiting

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

sincerely I'm confused, since the formula which models the temperature of a thermometer, is: R=R_0(1+At+Bt^2) where R_0, A , B are three coefficients. Nevertheless we have not enough data in order to determine those cefficients

OpenStudy (anonymous):

ok. no probs. what about this?

OpenStudy (anonymous):

The temperature at which the tungsten filament of a 12 V and 36W lamp operates is 1730o C. If the temperature coefficient of resistance of tungsten is 6×10−3 /K, find the resistance of the lamp at a room temperature of 20oC 10.00Ω 0.45Ω 0.39Ω 4.00Ω

OpenStudy (michele_laino):

here we have to apply the formual above, namely: \[R\left( t \right) = {R_0}\left( {1 + \alpha t} \right)\] where \alpha is 6×10−3 and t= 20°C. So, what do you get?

OpenStudy (anonymous):

R(t)=R(1.12)

OpenStudy (michele_laino):

and R_0= 1730 ohms so: R=1730*1.12=...?

OpenStudy (anonymous):

1937.6

OpenStudy (anonymous):

so, what next sir?

OpenStudy (michele_laino):

I got: 0.006*20=0.12

OpenStudy (anonymous):

ok. R=1730*0.12=207.6

OpenStudy (anonymous):

what next sir?

OpenStudy (michele_laino):

oops.. we have: R=4 ohm at t=1730 °C now we want to compute the resistance at t=20°C

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

please give the solution sir

OpenStudy (michele_laino):

we have to solve this equation: \[\frac{4}{x} = \frac{{1 + 0.006 \times 1730}}{{1 + 0.006 \times 20}}\] what is x?

OpenStudy (anonymous):

0.046